The product of two consecutive integers is 41 more than their sum. Find the integers
Answer by checkley77(12844)

x(x+1)=41+x+(x+1)
x^2+x=41+2x+1
x^2+x-2x-41-1=0
x^2-x-42=0
(x-7)(x+6)=0
x-7=0
x=7 ans.
7+1=8 ans.
x+6=0
x=-6 ans.
-6+1=-5 ans.
Proofs:
7*8=41+7+8
56=56
-6*-5=41-6-5
30=41-11
30=30